OpenStudy (vshiroky):

Averages... will medal! Accidentally started something and now I can't leave for work until I finish!!!

OpenStudy (phi):

this question is not really about averages. It is solving 2 equations 3x+2y= 79.05 2x+3y= 75.95

OpenStudy (phi):

set C=R and solve for x

ganeshie8 (ganeshie8):

solve below ``` t+s = 3889.5 t = s-689.5 ```